The base of the rectangle is changing at the rate of 3in/min. if its height remains constant, determine the rate of change of its perimeter with respect to time?


1
Expert's answer
2022-03-24T03:09:35-0400

"\frac{da}{dt}=3in/\min \\b=Const\\P=2a+2b\\\frac{dP}{dt}=2\frac{da}{dt}=6in/\min"